OpenStudy (anonymous):

Find the solutions to the equation below. 6x2+7x-5=0

OpenStudy (asylum15):

6x squared?

OpenStudy (anonymous):

its saying theres more than one solution but im having problems getting to that step

OpenStudy (asylum15):

6x\[6x ^{2}\]

OpenStudy (asylum15):

?

OpenStudy (anonymous):

yes its 6x ^2

OpenStudy (anonymous):

6x^(2)+7x+-5=0 (3x+5)(2x-1)=0 3x+5=0 2x-1=0 x=-5 / 3, 1 /2

OpenStudy (asylum15):

Damn you Luis! :D

OpenStudy (anonymous):

LOL..thanks alot you guys !!!

OpenStudy (anonymous):

kk